OpenStudy (anonymous):

Can someone help me solve this? \[(2+\sqrt{5}) (6-2\sqrt{5})^{2}\] Answer: a+b\sqrt{5}

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[(2+\sqrt{5}) (6-2\sqrt{5})^{2}\]

OpenStudy (anonymous):

Answer: \[a+b \sqrt{5}\]

OpenStudy (helder_edwin):

first expand \[ (6-2\sqrt{5})^2=... \]

OpenStudy (anonymous):

I got \[-28+28\sqrt{5}\] Did you get something else?

OpenStudy (helder_edwin):

u must know this \[\large (a-b)^2=a^2-2ab+b^2 \]

OpenStudy (helder_edwin):

so \[\large (6-2\sqrt{5})^2=6^2-2\times6\times2\sqrt{5}+[2\sqrt{5}]^2 =36-24\sqrt{5}+4\times5\] \[\large =36-24\sqrt{5}+20=56-24\sqrt{5} \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

\[56-24\sqrt{5}\] That can't be the answer tho.. Thanks anyway :)

OpenStudy (helder_edwin):

we haven't finished

OpenStudy (helder_edwin):

so \[\large (2+\sqrt{5})(6-2\sqrt{5})^2=(2+\sqrt{5})(56-24\sqrt{5})=... \] can u do this?